OpenStudy (urturn):

write equation for the line that is parrallel to given line and passes through the given point y=3x+7; (2,10)

OpenStudy (anonymous):

They give you y=3x+7 and (2,10)

OpenStudy (anonymous):

To get a line that is parallel, you keep the same slope but change the y-intercept.

OpenStudy (anonymous):

In this case all that is need from the function given is the slope.

OpenStudy (anonymous):

What is the slope in that function?

OpenStudy (urturn):

3 and 7

OpenStudy (urturn):

that the slope

OpenStudy (anonymous):

Why 3 and 7? Slope is equal to the coefficient of he x variable. So it is 3.

OpenStudy (urturn):

u want me show u the equation in a pic

OpenStudy (anonymous):

\[m=3\]\[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

No need for a pic.

OpenStudy (urturn):

ok

OpenStudy (anonymous):

The equation above is called point-slope form. It can help you derive an equation using just the coordinates of a point and the slope

OpenStudy (anonymous):

m= 3 x= 2 y=10 All you do is plug into that formula ans solve for y.

OpenStudy (urturn):

do want to see the answer they put up

OpenStudy (urturn):

y=33x+11 y=3x+4 y=-3x+4 y=(-1/3)x+4

OpenStudy (anonymous):

Which one is it?

OpenStudy (anonymous):

Follow my explanation and solve.

OpenStudy (urturn):

ok

OpenStudy (urturn):

help